# Optimizing Frequent Checkpointing via Low-Cost Differential for Distributed Training Systems

### v2

@@ Abstract
We proposes LowDiff, an efficient frequent checkpointing framework that reuses compressed gradients, serving as differential checkpoints to reduce cost. Furthermore, LowDiff incorporates a batched gradient write optimization to persist these differentials to storage efficiently. It also dynamically tunes both the checkpoint frequency and the batching size to maximize performance. In non-compression scenario, We further proposes LowDiff+ with a layer-wise gradient reusing and snapshotting approach and a CPU-based asynchronous persistence strategy, enabling frequent checkpointing without gradient compression. Experiments on various workloads show that LowDiff can achieve checkpointing frequency up to per iteration with less than 3.1% runtime overhead.
@@ I Introduction
Deep learning has demonstrated remarkable capabilities in the fields of computer vision [simonyan2014very, he2016deep], natural language processing [devlin2018bert, brown2020language] and so on. Recently, the advent of large models like ChatGPT [chatgpt] and GPT-4 [openai2024gpt4] has captivated both academic and industrial attention. To efficiently train the large models, distributed training systems [li2013parameter, li2014communication, sergeev2018horovod, rasley2020deepspeed] are widely deployed across workers (or nodes) to accelerate the training process, which iterates over a dataset many times (i.e., epochs), each containing multiple iterations. Nevertheless, distributed training often requires many nodes and a long training time, leading to frequent failures [zhang2020empirical]. For instance, a Microsoft study reveals that the mean time between failures (MTBF) in multi-tenant GPU clusters can range from minutes to days [jeon2019analysis]; similarly, OPT-175B, equipped with 992 NVIDIA A100 GPUs, experienced 112 failures during its 90-day training period—averaging nearly two failures per day [OPT].
@@
Therefore, to address the above challenges, the corresponding findings inspire us to leverage the compressed gradients in distributed training systems for low-cost differential checkpointing, so as to improve the overall training performance. Our motivation is that the compressed gradients can act as the compressed differentials with no DC computation cost and with a lower DC transmission cost. Based on the findings, we propose LowDiff, an efficient frequent checkpointing framework, whose main idea is to reuse the compressed gradients to serve as differential checkpoints, which not only eliminates the computation cost for differential compression (via reusing) but also reduces the transmission cost for DC writes (due to smaller-size gradients than differentials). Further, to enhance LowDiff, we accelerate the checkpoint writes via gradient batching and optimize the checkpointing performance by tuning the checkpoint frequency and batching size.
In scenarios without gradient compression, we further investigate how to efficiently reuse gradients for frequent DC construction and propose LowDiff+. In this case, directly compressing differentials would incur high computation costs, so we reuse and snapshot gradients as they are generated layer by layer during the backward pass, overlapping their offloading with computation, thus enabling a fine-grained reusing and snapshotting approach. Meanwhile, the model state is reconstructed in CPU memory by the reused gradients and regularly persisted to storage, both in parallel with ongoing GPU training, thus enabling high-frequency in-memory checkpointing and asynchronous persistence.
@@
  • We perform an analysis of LowDiff, showing that gradient reuse eliminates the data dependency and enables parallel execution between checkpointing and training. Through measurement, we further demonstrate that checkpointing can be overlapped with the training iteration, allowing checkpointing to execute concurrently with training (§IV).
@@
  • We further propose LowDiff+ for efficient gradient reuse without compression. By exploiting the layer-by-layer generation of gradients during training, we design a layer-wise gradient reusing and snapshotting approach. We also reconstruct a model state in CPU memory, performing high-frequency in-memory checkpointing and asynchronous persistence (§VI).
  • We implement LowDiff and LowDiff+ atop DeepSpeed, open-sourced at https://github.com/YuchongHu/LowDiff. Experiments on various models show that LowDiff and LowDiff+ can achieve checkpointing as frequently as once per iteration with less than minimal runtime overhead. Compared to the state-of-the-art methods, LowDiff and LowDiff+ can shorten training time by up to $89.2\%$ and $81.7\%$ under high checkpointing frequencies (§VII & §VIII).
@@ II-B Checkpointing Techniques
Then the model state $M_{t}$ at the $t$-th iteration can be reconstructed by merging the full and all differential checkpoints as
@@ IV Analysis
By incorporating the main idea of gradient reuse (§III), LowDiff aims to enable the DC process to run concurrently with model training, thereby minimizing potential stalls. To confirm the feasibility of such parallelism, we analyze our proposed strategy from two key aspects:
  • Parallelism property: it ensures no data dependencies block concurrency with training iterations.
  • Execution time overlap: it reveals whether the checkpointing cost can be effectively hidden within the training time.
@@
Figure 3: LowDiff’s DC can run in parallel with forward pass(F), backward pass(B), and model update(U).
IV-A Parallelism Analysis
@@
Necessity of Parallelism Analysis. Checkpointing can be executed with the forward and backward passes which do not modify the model state [mohan2021checkfreq], but it cannot run in parallel with model update which directly writes to it [wang2024fastpersist]. In other words, checkpointing is data-dependent on the model update [wang2024fastpersist]. The reason is that it must read the model state after model updates (Read-after-Write [hennessy2011computer], RAW dependency), and the model update at the next iteration must write the model state after checkpointing completes its read (Write-after-Read [hennessy2011computer], WAR dependency). Consequently, the two dependency patterns incur training stalls, preventing from paralleling checkpointing with model update. Prior studies [mohan2021checkfreq, wang2023gemini, wang2024fastpersist] design their checkpointing strategies based on these data dependencies to pipeline training and checkpointing.
@@
Consequently, it is necessary to carefully analyze: (1) the data dependency introduced by our reuse strategy and (2) the parallelism property of LowDiff’s DC, specified as follows.
@@
(1) LowDiff’s DC can eliminate the WAR dependency. We next analyze current DC’s WAR dependency with next iteration’s model update, backward, and forward pass.
@@
  • Model update: as described in §II-A, the model update (Equation (4)) only requires read-only access to the gradient. Because LowDiff checkpoints only the compressed gradient, leveraging gradient reuse, the original Write-after-Read (WAR) dependency is eliminated because the checkpointing process no longer needs to read the model state ($M_{t}$). Instead, it just reads the compressed gradient ($\widetilde{G}_{t}$), which is available in the backward pass and will not be changed after. Since both the checkpointing and model update are read-only consumers of the gradient, they can operate in parallel without conflict.
@@
  • Backward pass: LowDiff’s gradient reuse strategy introduces no additional WAR dependency on backward pass, since each gradient ${G}_{t}$ is finalized within its iteration and never modified thereafter. Specifically, once a gradient ${G}_{t}$ is generated in the current iteration ($t$), it is final and will not be modified by subsequent steps. The backward pass of the $t+1$-th iteration (Equation (2)) does not read or depend on ${G}_{t}$; it generates a completely new gradient ($G_{t+1}$) from scratch and won’t change the ${G}_{t}$. Therefore, the DC’s read of ${G}_{t}$ does not conflict with the $B$’s write of ${G}_{t+1}$ at $t+1$-th iteration, allowing LowDiff’s DC to run in parallel with the next iteration’s backward pass.
@@
  • Forward pass: The forward pass (Equation (1)) doesn’t involve gradient, so there is no WAR dependency between LowDiff’s DC and forward pass.
@@
(2) LowDiff can run in parallel with forward pass, backward pass, and model update. To visualize the parallelism benefit from data dependency, Figure 3 shows the execution timeline with LowDiff. We see that the DC task ($C^{D}_{t}$) at the $t$-th iteration is no longer coupled with the model update ($U$) or the following iteration (no dependency arrow from DC to training). Instead, the reusing operation on compressed gradient ($\widetilde{G}_{t}$) begins immediately after the backward pass ($B$), running in parallel with the model update and subsequent iteration’s forward pass(F), backward pass (B), and model update(U). This parallelism property effectively transforms LowDiff’s DC from a blocking operation to a parallel execution with training, thereby eliminating the synchronization stalls previously caused by the WAR dependency.
@@ IV-B Overlap Analysis
Necessity of Overlap Analysis. Beyond its parallelism property, LowDiff can perform even better if the checkpointing process is fast enough to be hidden by the training iteration, thus minimizing DC’s cost. Specifically, the time required to checkpoint a compressed gradient (denoted as $C$) should be shorter than a single training iteration (denoted as $F+B+U$), similar to the Microsoft study [mohan2021checkfreq].
Consequently, it is also necessary to analyze the potential of overlapping, so we first (1) take a measurement analysis of training time, full checkpointing, and DC time, as well as (2) discuss why the overlap can be practically feasible.
(1) The checkpointing of compressed gradients can be overlapped with training. Following the measurement setup in §III-A, we measure the iteration time, full checkpointing time and differential checkpointing time on BERT-B, BERT-L, GPT2-S, and GPT2-L. As shown in Figure 4, the DC time is only 20.5%, 22.1%, 23.1%, and 24.6% of the corresponding training iteration time for these models. This indicates that in most cases, the checkpointing process can be fully overlapped with training iterations without introducing noticeable stalls.
@@
Based on the above analysis, we confirm that the concurrent execution between LowDiff’s DC and training is theoretically feasible. While this analysis establishes the feasibility of frequent DC, the detailed design of LowDiff will be presented in §V.
@@ V Design
Figure 5 shows the architecture of LowDiff. Each worker has two processes: Training and Checkpointing, connected via Reusing Queue for gradient reuse. Training first generates the local gradient, compresses it in Compressor, and synchronizes it across workers via Communicator. The synchronized compressed gradient $\widetilde{G}_{t}$ is put into Reusing Queue via zero-copy technique for reuse in checkpointing and is also decompressed by Decompressor to update the model state. Training checkpoints the model state $M_{t}$ regularly as full checkpoints $C^{F}_{t}$. Checkpointing snapshots compressed gradients from Reusing Queue to form differential checkpoints $C^{D}_{t}$, which Batched Writing module caches in CPU memory and aggregates into a single batched differential checkpoint $C^{B}_{t}$. Both the full checkpoint frequency and batching size are dynamically tuned by Optimal Configuration module. Finally, the batched differential and full checkpoints are persisted to local or remote storage.
@@ V-A Reusing Compressed Gradients
Reusing with zero-copy optimization. To meet Requirement 2, we use the CUDA Inter-Process Communication (IPC) queue in PyTorch to implement a zero-copy mechanism. This allows GPU memory to be shared between training and checkpointing without extra copying, significantly reducing communication overhead. Specifically, the queue serves as a high-level data structure for reusing compressed gradients, but only transmits the memory handle. It relies on the CUDA IPC library to share the gradient’s GPU memory handle across processes, avoiding additional I/O overhead from memory copying. In PyTorch, this can be achieved with torch.multiprocessing.Queue(), where the inter-process communication creates an IPC handle when a tensor is enqueued and reopens it when dequeued.
Algorithm 1 Reusing in LowDiff
@@
4:    $G_{i,t}\leftarrow Backward(x_{t})$
@@
11:   while $\widetilde{G}_{t}\leftarrow{\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.get}}$ do $\triangleright$ Reusing
12:    $C^{D}_{t}\leftarrow Save(\widetilde{G}_{t})$ $\triangleright$ Diff CKPT
13:   $C^{F}_{t}\leftarrow Save(M_{t})$ $\triangleright$ Full CKPT Regularly
14: function Recovery Process
15:   $M_{t}\leftarrow Load(C^{F}_{t})$ $\triangleright$ Load Full CKPT
16:   for $j\leftarrow t$, …, $n$ do $\triangleright$ Restore to Latest
17:    $\widetilde{G}_{j}\leftarrow Load(C^{D}_{j})$ $\triangleright$ Load Diff CKPT
18:    $G_{j}\leftarrow Comp^{-1}(\widetilde{G}_{j})$ $\triangleright$ Decompress
19:    $M_{j+1}\leftarrow M_{j}+Adam(G_{j})$ $\triangleright$ Diff Merge
20:   return $M_{n}$
Design details Algorithm 1 specifies the following details of reusing compressed gradients, including training, checkpointing, and recovery processes.
Training process. After performing a backward pass and generating the gradient $G_{i,t}$ of worker $i$ at the $t$-th iteration (Lines 1-3), the training process compresses $G_{i,t}$ into $\widetilde{G}_{i,t}$ for communication (Line 4). Then, the compressed gradient is synchronized across all workers using a collective communication primitive (e.g., Allgather or Allreduce) (Line 5). Following synchronization, the training process puts the compressed gradient $\widetilde{G}_{t}$ into Reusing Queue $Q$ (Line 6), which will be transmitted to the checkpointing process with zero-copy. Finally, the training process decompresses the gradient (Line 6), updates the model state based on the gradient (Line 8), and proceeds to the next iteration.
Checkpointing process. The checkpointing process repeatedly retrieves the compressed gradients from Reusing Queue $Q$ to CPU memory (Line 11). Subsequently, the checkpointing process determines whether to merge the compressed gradient into batched differential checkpoints (detailed in §V-B). After that, it invokes torch.save to persist the compressed differential checkpoints $C^{D}_{t}$ (Lines 12-14) to storage. The model state is also regularly checkpointed as a full checkpoint $C^{F}_{t}$ at the $t$-iteration (Line 15).
Recovery process. During recovery, the recovery process first loads the latest full checkpoint $C^{F}_{t}$ at the $t$-th iteration to restore the model state $M_{t}$ (Line 17). Then, the compressed gradient $\widetilde{G}_{j}$ is retrieved from the differential checkpoint $C^{D}_{j}$ at the $j$-th iteration (Line 20) and decompressed into gradient $G_{j}$ (Line 21). Subsequently, the model state is restored from $M_{j}$ to $M_{j+1}$ at the $(j)$-th iteration by the optimizer and gradient (Line 21). By repeating this process for $n$ iterations (Lines 18-23), where compressed gradients from differential checkpoints are successively loaded and applied, the model state is finally restored to $M_{n}$ (Line 24), thus recovering the training from failures.

### v3

@@ Abstract
We propose LowDiff, an efficient frequent checkpointing framework that reuses compressed gradients, serving as differential checkpoints to reduce cost. Furthermore, LowDiff incorporates a batched gradient write optimization to persist these differentials to storage efficiently. It also dynamically tunes both the checkpoint frequency and the batching size to maximize performance. To enhance LowDiff under non-compression scenarios, we further propose LowDiff+, which incorporates a layer-wise-reuse snapshotting strategy, along with an incremental-merging persistence strategy. Experiments on various workloads show that LowDiff and LowDiff+ can reduce the training time by up to 89.2% and 81.2% with checkpointing frequency up to per iteration.
@@ I Introduction
Deep learning has demonstrated remarkable capabilities in the fields of computer vision [simonyan2014very, he2016deep], natural language processing [devlin2018bert, brown2020language] and so on. Recently, the advent of large models like ChatGPT [chatgpt] and GPT-4 [openai2024gpt4] has captivated both academic and industrial attention. To efficiently train the large models, distributed training systems [li2013parameter, li2014communication, sergeev2018horovod, rasley2020deepspeed] are widely deployed across workers (or nodes) to accelerate the training process, which iterates over a dataset many times (i.e., epochs), each containing multiple iterations. Nevertheless, distributed training often requires many nodes and a long training time, leading to frequent failures [zhang2020empirical]. For instance, a Microsoft study reveals that the mean time between failures (MTBF) in multi-tenant GPU clusters can range from minutes to days [jeon2019analysis]; similarly, OPT-175B, equipped with 992 NVIDIA A100 GPUs, experienced 112 failures during a 90-day training period, averaging nearly two failures per day [OPT].
@@
Therefore, to address the above challenges, the corresponding findings inspire us to leverage the compressed gradients in distributed training systems for low-cost differential checkpointing, so as to improve the overall training performance. Our motivation is that the compressed gradients can act as the compressed differentials with no DC computation cost and with a lower DC transmission cost. Based on the findings, we propose LowDiff, an efficient frequent checkpointing framework, whose main idea is to reuse the compressed gradients to serve as differential checkpoints, which not only eliminates the computation cost for differential compression (via reusing) but also reduces the transmission cost for DC writes (due to smaller-size gradients than differentials). Further, we accelerate the checkpoint writes via gradient batching and optimize the checkpointing performance by tuning the checkpoint frequency and batching size.
We further investigate how to reuse gradients for frequent checkpointing in scenarios without gradient compression. We first identify the issues of snapshotting and persistence when dealing with uncompressed gradients. To address these, we offer two key insights: (1) gradients can be reused layer-by-layer, and (2) differential and full checkpoints can be merged together. Based on the two insights, we proposed LowDiff+, which employs a layer-wise-reuse snapshotting strategy, along with an incremental-merging persistence strategy, thus enabling high-frequency checkpointing without gradient compression.
@@
  • We perform an analysis of LowDiff, showing that gradient reuse eliminates the data dependency and enables parallel execution between checkpointing and training. Through measurement, we further demonstrate that checkpointing time can be overlapped with the training iteration, allowing checkpointing to execute concurrently with training (§IV).
@@
  • We further propose LowDiff+ for efficient gradient reuse without compression. By exploiting the layer-by-layer generation of gradients during training, LowDiff+ reuses and snapshots gradient in layer-wise granularity. LowDiff+ also merge the full model state and differentials in CPU memory before persistence to eliminates the overhead of writing separate differentials to storage (§VI).
  • We implement LowDiff and LowDiff+ atop DeepSpeed, open-sourced at https://github.com/YuchongHu/LowDiff. Experiments on various models show that LowDiff and LowDiff+ can achieve checkpointing as frequently as once per iteration with minimal runtime overhead. Compared to the state-of-the-art methods, LowDiff and LowDiff+ can shorten training time by up to $89.2\%$ and $81.7\%$ under high checkpointing frequencies (§VII & §VIII).
@@ II-B Checkpointing Techniques
Then the model state $M_{t}$ at the $t$-th iteration can be reconstructed by merging the full and differential checkpoints as
@@ IV Analysis
By incorporating the main idea of gradient reuse (§III), LowDiff aims to enable the checkpointing process to run concurrently with model training, thereby minimizing potential stalls. To confirm the feasibility of such parallelism, we analyze our proposed strategy from two key aspects:
  • Parallel of Checkpointing: it ensures no data dependencies block concurrency with training iterations.
  • Overlap of execution time: it reveals whether the checkpointing cost can be effectively hidden within the training time.
@@
Figure 3: LowDiff can run in parallel with forward pass ($F$), backward pass ($B$), and model update ($U$).
IV-A Parallel of Checkpointing
@@
Necessity of Parallelism Analysis. Checkpointing can be executed with the forward and backward passes, which do not modify the model state [mohan2021checkfreq], but it cannot run in parallel with model update, which directly writes to it [wang2024fastpersist]. In other words, checkpointing is data-dependent on the model update [wang2024fastpersist]. The reason is that it must read the model state after model updates (Read-after-Write [hennessy2011computer], RAW dependency), and the model update at the next iteration must write the model state after checkpointing completes its read (Write-after-Read [hennessy2011computer], WAR dependency). Consequently, the two dependency patterns incur training stalls, preventing parallelism of checkpointing with model update. To visualize such data dependency, Figure 3(a) shows the execution timeline with checkpointing. We see that the first dependency on $M_{t}$ (shown by the arrows from $U$ to $C$) and the second dependency on $M_{t}$ (shown by the arrows from $C$ to $U$) create a stall ($S$) before model update $U$ of the $i+1$-th iteration. Prior studies [mohan2021checkfreq, wang2023gemini, wang2024fastpersist] design their checkpointing strategies based on these data dependencies to pipeline training and checkpointing.
@@
Consequently, it is necessary to carefully analyze the data dependency introduced by our reuse strategy and the parallelism of checkpointing in LowDiff, specified as follows.
@@
LowDiff can eliminate the WAR dependency. We next analyze the WAR dependency with the next iteration’s model update, backward, and forward pass.
@@
  • Model update. As described in §II-A, the model update (Equation (4)) only requires read-only access to the gradient. Because LowDiff checkpoints only the compressed gradient, leveraging gradient reuse, the original Write-after-Read (WAR) dependency is eliminated because the checkpointing process no longer needs to read the model state ($M_{t}$). Instead, it just reads the compressed gradient ($\widetilde{G}_{t}$), which is available in the backward pass and will not be changed after. Since both the checkpointing and model update are read-only consumers of the gradient, they can operate in parallel without conflict.
@@
  • Backward pass. LowDiff’s gradient reuse strategy introduces no additional WAR dependency on backward pass, since each gradient ${G}_{t}$ is finalized within its iteration and never modified thereafter. Specifically, once a gradient ${G}_{t}$ is generated in the current iteration ($t$), it is final and will not be modified by subsequent steps. The backward pass of the $(t+1)$-th iteration (Equation (2)) does not read or depend on ${G}_{t}$; it generates a completely new gradient $G_{t+1}$ from scratch and won’t change ${G}_{t}$. Therefore, the DC’s read of ${G}_{t}$ does not conflict with the backward pass’s write of ${G}_{t+1}$ at the $(t+1)$-th iteration, allowing LowDiff to run in parallel with the next iteration’s backward pass.
@@
  • Forward pass. The forward pass (Equation (1)) doesn’t involve the gradient, so there is no WAR dependency between LowDiff and the forward pass.
@@
LowDiff can run in parallel with forward pass, backward pass, and model update. To visualize the parallelism benefit from data dependency, Figure 3(b) shows the execution timeline with LowDiff. We see that the checkpointing ($C_{t}$) at the $t$-th iteration is no longer coupled with the model update ($U$) or the following iteration (no dependency arrow from checkpointing to training). Instead, the reusing operation on compressed gradient ($\widetilde{G}_{t}$) begins immediately after the backward pass ($B$), running in parallel with the model update and subsequent iteration’s forward pass ($F$), backward pass ($B$), and model update ($U$). This parallelism property effectively transforms LowDiff from a blocking operation to a parallel execution with training, thereby eliminating the synchronization stalls previously caused by the WAR dependency.
@@ IV-B Overlap Analysis
Necessity of Overlap Analysis. Beyond its parallelism property, LowDiff can perform even better if the checkpointing process is fast enough to be hidden by the training iteration, thus minimizing DC’s cost. In other words, the time required to checkpoint a compressed gradient (denoted as $C$) should be shorter than a single training iteration (denoted as $F+B+U$), similar to the Microsoft study [mohan2021checkfreq].
Consequently, it is also necessary to analyze the potential of overlapping, so we first (1) conduct a measurement analysis of training time, full checkpointing, and DC time, as well as (2) discuss why the overlap can be practically feasible.
(1) The checkpointing of compressed gradients can be overlapped with training. Following the measurement setup in §III-A, we measure the iteration time, full checkpointing time, and differential checkpointing time on BERT-B, BERT-L, GPT2-S, and GPT2-L. As shown in Figure 4, the DC time is only 20.5%, 22.1%, 23.1%, and 24.6% of the corresponding training iteration time for these models. This indicates that in most cases, the checkpointing process can be fully overlapped with training iterations without introducing noticeable stalls.
@@
Based on the above analysis, we confirm that the overlap of concurrent executions of LowDiff’s DC and training is practically feasible.
@@ V Design
Figure 5 shows the architecture of LowDiff. Each worker has two processes: Training and Checkpointing, connected via Reusing Queue for gradient reuse. Training first generates the local gradient, compresses it in Compressor, and synchronizes it across workers via Communicator. The synchronized compressed gradient $\widetilde{G}_{t}$ is enqueued into Reusing Queue via zero-copy technique for reuse in checkpointing and is also decompressed by Decompressor to update the model state. Training checkpoints the model state $M_{t}$ regularly as full checkpoints $C^{F}_{t}$. Checkpointing snapshots compressed gradients from Reusing Queue to form differential checkpoints $C^{D}_{t}$, which Batched Writing module caches in CPU memory and aggregates into a single batched differential checkpoint $C^{B}_{t}$. Both the full checkpoint frequency and batching size are dynamically tuned by Optimal Configuration module. Finally, the batched differential and full checkpoints are persisted to local or remote storage.
@@ V-A Reusing Compressed Gradients
Reusing with zero-copy optimization. To meet Requirement 2, we use the CUDA Inter-Process Communication (CUDA IPC) queue in PyTorch to implement a zero-copy mechanism. This allows GPU memory to be shared between training and checkpointing without extra copying, significantly reducing communication overhead. Specifically, the queue serves as a high-level data structure for reusing compressed gradients, but only transmits the memory handle. It relies on the CUDA IPC library to share the gradient’s GPU memory handle across processes, avoiding additional I/O overhead from memory copying. In PyTorch, this can be achieved with torch.multiprocessing.Queue(), where the inter-process communication creates an IPC handle when a tensor is enqueued and reopens it when dequeued.
Algorithm 1 LowDiff
@@
4:    $G_{i,t}\leftarrow Backward(x_{t})$ $\triangleright$ Backward Pass
@@
11:   for all $t$ do
12:    while $\widetilde{G}_{t}\leftarrow{\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.get}}$ do $\triangleright$ Reusing
13:      $C^{D}_{t}\leftarrow Save(\widetilde{G}_{t})$ $\triangleright$ Diff CKPT
14:    $C^{F}_{t}\leftarrow Save(M_{t})$ $\triangleright$ Full CKPT Regularly
15: function Recovery Process
16:   $M_{t}\leftarrow Load(C^{F}_{t})$ $\triangleright$ Load Full CKPT
17:   for $j\leftarrow t$, …, $n$ do $\triangleright$ Restore to Latest
18:    $\widetilde{G}_{j}\leftarrow Load(C^{D}_{j})$ $\triangleright$ Load Diff CKPT
19:    $G_{j}\leftarrow Comp^{-1}(\widetilde{G}_{j})$ $\triangleright$ Decompress
20:    $M_{j+1}\leftarrow M_{j}+Adam(G_{j})$ $\triangleright$ Diff Merge
21:   return $M_{n}$
Design details. Algorithm 1 specifies the following details of reusing compressed gradients, including training, checkpointing, and recovery processes.
Training process. After performing a backward pass and generating the gradient $G_{i,t}$ of worker $i$ at the $t$-th iteration (Lines 1-3), the training process compresses $G_{i,t}$ into $\widetilde{G}_{i,t}$ for communication (Line 4). Then, the compressed gradient is synchronized across all workers using a collective communication primitive (e.g., Allgather or Allreduce) (Line 5). Following synchronization, the training process puts the compressed gradient $\widetilde{G}_{t}$ into Reusing Queue $Q$ (Line 6), which will be transmitted to the checkpointing process with zero-copy. Finally, the training process decompresses the gradient (Line 7), updates the model state based on the gradient (Line 8), and proceeds to the next iteration.
Checkpointing process. The checkpointing process repeatedly retrieves the compressed gradients from Reusing Queue $Q$ to CPU memory (Line 11). Subsequently, the checkpointing process determines whether to merge the compressed gradient into batched differential checkpoints (detailed in §V-B). After that, it invokes torch.save to persist the compressed differential checkpoints $C^{D}_{t}$ (Lines 12) to storage. The model state is also regularly checkpointed as a full checkpoint $C^{F}_{t}$ at the $t$-iteration (Line 13).
Recovery process. During recovery, the recovery process first loads the latest full checkpoint $C^{F}_{t}$ at the $t$-th iteration to restore the model state $M_{t}$ (Line 15). Then, the compressed gradient $\widetilde{G}_{j}$ is retrieved from the differential checkpoint $C^{D}_{j}$ at the $j$-th iteration (Line 17) and decompressed into gradient $G_{j}$ (Line 18). Subsequently, the model state is restored from $M_{j}$ to $M_{j+1}$ at the $(j)$-th iteration by the optimizer and gradient (Line 19). By repeating this process for $n$ iterations (Lines 16), where compressed gradients from differential checkpoints are successively loaded and applied, the model state is finally restored to $M_{n}$ (Line 20), thus recovering the training from failures.